Incorrect Question 8 0/1 pts 0.28 litre of an ideal monatomic gas (Cv.m 3R/2) initially at 2 °C and 95 atm pressure undergo an expansion against a constant external pressure of 1.02 atm, and do 1.2 k of work. The final pressure of the gas is 1.02 atm. Calculate the change in enthalpy, AH. Report your answer in J 3660
Answer #1

Work done=pex.(∆V)=1.2 kJ=1200 J=1.02atm( Vf-0.28)

1200=1.02×101.33J(Vf-0.28)

Vf=11.89 L

Tf=Vf.Pf/ R=11.89×1.02/0.082=147.9 K

Ti=2+273=275 K.

∆T=147.9-275=-127.1 K

Cp=Cv+R=5/2 R

∆H=nCp.∆T

=5/2R×-127.1

=-2641.8 J